OpenStudy (anonymous):

if three times a number added to 2 is divided by the number plus 5, the result is five halves, find the number

OpenStudy (anonymous):

Let x=the number. Your question says: \[\frac{ 3x+2 }{ x+5 } = \frac{ 5 }{ 2 }\]And now you need to solve for x. You can start by cross multiplying.